# Diffusion Language Models Are Versatile Protein Learners

### v1

@@ B.1 Pre-training of DPLM
During the training phase, we investigate two different approaches: (1) training with the diffusion objective from scratch, and (2) what we refer to as two-stage training, which consists of initial training with the masked language modeling (MLM) objective followed by continuous training with the diffusion objective. We find that training DPLM from scratch is a bit challenging and converges slowly. Training with MLM first allows the model to get a better starting point, and then get the final model through diffusion adaptation, sharing similar idea as Ye et al. (2023a). This kind of strategy similar to curriculum learning makes the training effect better. We can also use community available pre-trained Masked-LMs such as ESM2 (Lin et al., 2022) as our first-stage model. We found that the validation loss of two-stage training from our in-house trained Masked-LM and pre-trained ESM2 is very close, and as shown in Fig, there is also almost no performance gap in the unconditional sampling.
B.2 Ar-LM baseline
@@
We pretrain a Ar-LM using autoregressive training objective. In order to be comparable with DPLM, the autoregressive language model we trained adopts the same architecture as DPLM, which is the same architecture as ESM2. To be capable of adapting the autoregressive training, we modify the mask matrix of the attention module to causal mask, which guarantees each token can only attend the previous position and keep unseen for future. The training objective is next word prediction, and we process the input sequence with teacher forcing for efficient parallel training. During decoding, we start with <bos> token, sample one token each timestep from left to right, and the sampled token in the current timestep will be concatenated to the end of the sequence, becoming input for next timestep. The decoding process terminates until <eos> token is sampled. Because we can not know when to obtain the <eos> token in advance, we can not decide the length of sampled sequence. We attempt to force the sampling length by modifying the sampling probability: the probability of <eos> is 1 when and only when the sequence length is up to the predefined length, while 0 in all the previous timesteps. However, we observe this will decline the quality of sampled sequence significantly.
B.3 A modified unconditional sampling strategy
@@
Appendix C Additional Experimental Results
C.1 Sequence-conditional generation: motif-scaffolding
@@
C.2 Structure-conditional generation: inverse folding
@@
$\mathcal{J}_{t}=\mathbb{E}_{q(\mathbf{x}^{(0)})}\bigg{[}\lambda^{(t)}\sum_{1\leq i\leq L}b_{i}(t)\cdot\log p_{\theta}(\mathbf{x}^{(0)}_{i}|\mathbf{x}^{(t)}_{draft})\bigg{]},$ (7)
@@
C.3 Classifier-free guidance
@@
Appendix D Related Work
D.1 Language Models
The dominant paradigm of language models is autoregressive language models, which breaks down the mutual distribution over the tokens of a sequence into conditional probabilities via the chain rule ${p(\mathbf{x}^{[1:N]})=\prod_{i=1}^{N}p(x^{[i]}|\mathbf{x}^{[1:i-1]})}$ and generates tokens by ancestral sampling from left to right (Bengio et al., 2000; Sutskever et al., 2014; Vaswani et al., 2017). Recently, researchers propose the non-autoregressive language models as an alternative (Gu et al., 2018). These models do not need to obey the left to right generation order (Qian et al., 2022; Huang et al., 2023) and demonstrate competitive or superior performance compared to their autoregressive counterpart across a wide range of domains including languages (Qian et al., 2021; Huang et al., 2023; Qian et al., 2022; Huang et al., 2023; Zheng et al., 2023a), speeches (Kim et al., 2021), proteins (Zheng et al., 2023b), and molecules (Hoogeboom et al., 2022). Among the numerous non-autoregressive language models, diffusion language models (Li et al., 2022; Gong et al., 2022; Zheng et al., 2023a) have emerged as a solid and promising framework. Pretraining language models on a massive scale of unlabeled data markedly improves their downstream task performance (Mikolov et al., 2013; Peters et al., 2018; Radford et al., 2018; Devlin et al., 2019). As data volume and model sizes scale up, the training loss of language models predictably declines (Kaplan et al., 2020; Hoffmann et al., 2022; Muennighoff et al., 2023), and enhancing downstream task performance even without specific tuning (Radford et al., 2019). GPT3 (Brown et al., 2020) is a significant point in the journey, taking model sizes to 175B parameters, proposing in-context learning to bolster language models’ competence in solving certain tasks with only a handful of demonstrations. Furthermore, Wei et al. (2021); Sanh et al. (2022); Ouyang et al. (2022) introduce instruction tuning, finetuning pretrained language models on series of tasks described via instructions, which elicits the instruction following ability of models and significantly enhances their zero-shot performance on unseen tasks. More impressively, sufficiently large language models exhibit the emergent abilities such as multi-step reasoning (Kojima et al., 2022; Wei et al., 2022a, b), which small models do not possess (Fu et al., 2023). Empowered by large language models, helpful applications such as conversational AI systems [^1] https://chat.openai.com/ and autonomous agents [^2] https://github.com/Significant-Gravitas/Auto-GPT have garnered much interest. Although the most capable models at the moment are restricted in access, open-sourced efforts (Zeng et al., 2022; Touvron et al., 2023a, b; Taori et al., 2023; Chiang et al., 2023; Sun & Qiu, 2023) have largely enhanced the public accessibility of powerful large language models.
D.2 Protein Language Models
@@
D.3 Diffusion Language Models
@@
D.4 Protein Structure Diffusion Models
@@
D.5 Protein Inverse Folding
@@
Appendix E Visualization of Unconditional Samples

### v2

@@ B.1 Pre-training of DPLM
During the training phase, we investigate two different approaches: (1) training with the diffusion objective from scratch, and (2) what we refer to as two-stage training, which consists of initial training with the masked language modeling (MLM) objective followed by continuous training with the diffusion objective.
Empirical Observation.
In our preliminary experiments, we observed that discrete diffusion pre-training ”from scratch (FS)” often yielded instability in the form of frequent loss spiking, therefore hurting our model performance.
Our Hypothesis.
We noticed that the absorbing diffusion objective leads to a variable masking ratio ranging from 0% to 100%, while conventional MLM objective’s masking ratio keep fixed at 15% such that a masked LM is always exposed to rich condition of 85% observation/context. In other words, in contrast to MLM, for absorbing discrete diffusion models like DPLM, in some of the extreme cases there are nearly all tokens getting masked, which means that the model is required to recover all tokens of the ground-truth sequence from nothing). This could impose severe learning challenges, especially at the early phase of pre-training, where the model has yet not acquired sufficient capability to extract informative features and correlations from limited observation.
Solution in Principle.
Inspired by the success of curriculum learning in non-autoregressive text generation (Qian et al., 2021a; Gu & Kong, 2021; Guo et al., 2020a, b; Wang et al., 2022), we suggested that a masking warmup strategy could mitigate this issue, where we can start with a small upper bound of masking ratio (e.g., 15% as conventional MLM, to preserve a high proportion of observation) in the early phase of pre-training, and then gradually increase the masking ratio towards the authentic discrete diffusion objective during pre-training.
Solution in Practice.
In the current form of our manuscript, adhering to this principle, we proposed a two-stage training method: initialized DPLM from an established masked LM, either from our in-house pre-trained one or from the official ESM-2 checkpoint, and then trained DPLM by the discrete diffusion language modeling objective afterwards. Though it may or may not lead to the best model performance, it offers the possibilities of standing on the shoulders of any pre-trained masked LM such as ESM2 or the advanced LM architecture such as a Llama/Mistral-style masked LMs, in the broad open-source AI community. This also enables us to bypass the time-consuming process of gradual masking warmup during pre-training. As a result, this can be the most efficient and effective approach in practice, which also shares a similar principle as finetuning from RosettaFold in RFDiffusion (Watson et al., 2023).
Although discrete diffusion pre-training from scratch is challenging, we have also explored several ways to improve it. As shown in the Fig. 2G, DPLM can achieve comparable performance through pre-training from scratch. More specifically, we found that (1) gradient norm clipping can effectively help stabilize training process of discrete diffusion language modeling, greatly reducing the chance of loss spiking and gradient nan. In addition, we also found that (2) training longer (i.e. scaling compute) is another key to attain a good ultimate model performance (trained for 300k steps).
We are also curious about the performance of the vanilla masking warmup strategy and would like to see if it can lead to a better pre-trained DPLM. The training dynamics of discrete diffusion based DPLM is an interesting and exciting direction deserving further exploration, and we leave these as our future work.
B.2 Pre-training of Ar-LM baseline
@@
We pretrain a Ar-LM using autoregressive training objective. In order to be comparable with DPLM, the autoregressive language model we trained adopts the same architecture as DPLM. To be capable of adapting the autoregressive training, we modify the mask matrix of the attention module to causal mask, which guarantees each token can only attend the previous position and keep unseen for future. The training objective is next word prediction, and we process the input sequence with teacher forcing for efficient parallel training. During decoding, we start with <bos> token, sample one token each timestep from left to right, and the sampled token in the current timestep will be concatenated to the end of the sequence, becoming input for next timestep. The decoding process terminates until <eos> token is sampled. Because we can not know when to obtain the <eos> token in advance, we can not decide the length of sampled sequence. We attempt to force the sampling length by modifying the sampling probability: the probability of <eos> is 1 when and only when the sequence length is up to the predefined length, while 0 in all the previous timesteps. However, we observe this will decline the quality of sampled sequence significantly.
Appendix C Reasons for choosing absorbing discrete diffusion
We employed absorbing discrete diffusion as the pre-training method, instead of other forms of discrete diffusion or latent diffusion, for the following considerations.
C.1 Regarding discrete diffusion (DD) definitions (multinomial vs absorbing as proposed in D3PM)
Intuition.
We favor absorbing DD since the learning objective of absorbing DD generalizes existing language modeling objectives, as highlighted in section 4 of Austin et al. (2021), In particular, absorbing DD is a natural extension of the masked language modeling (MLM) objective, which has been thoroughly studied (Wettig et al., 2023) in the field of NLP and widely proven to be a robust and effective sequence learning protocol. In contrast, multinomial/uniform-DD resembles (tranfitional) denoising autoencoders (Savinov et al., 2021). There is little solid evidence and remains highly unclear about (mutlinomial) denoising autoencoder as a sequence learning objective can scale up w.r.t data, model size and application scenarios.
Empirical verification.
In our preliminary exploration, we have studied the performance of multinomial/uniform-DD and absorbing-DD. Here we provide the result regarding unconditional sampling, as shown in Fig. 5. We can find that DPLM-absorbing generally manifests better performance than DPLM-multinomial across different lengths.
Figure 5: The unconditional sampling performance of multinomial/uniform-DD and absorbing-DD.
C.2 Regarding latent diffusion
We reason that latent diffusion for discrete sequence data, which performs Gaussian diffusion in continuous embedding space, requires additional lossy continuous relaxation of discrete sequence data like protein sequence, which does not fit the discrete nature of protein sequence and is not necessarily the best choice for modeling discrete sequence data. Recent study (Meshchaninov et al., 2024) presents a latent diffusion model on protein LM embedding, where the pLDDT score unconditional sampling attains reasonable pLDDT in their paper, while our discrete diffusion based approach still excels.
Appendix D Additional Experimental Details
D.1 A modified unconditional sampling strategy
@@
D.2 Delve deeply into the pLDDT score of unconditional sampling
According to the Fig. 2A, we surprisingly find that the pLDDT score of DPLM unconditional sampling is even higher than the UniRef50 dataset. We investigate this phenomenon as follows.
Regarding the lower pLDDT in UniRef50.
The lower pLDDT here is because UniRef50 contains some data with lower structural plausibility, such as sequences with a large number of repetitive patterns. These cases decrease the average pLDDT of UniRef50, for instance sequence ADADAD...ADADAD with pLDDT 35.14.
We also investigate the average pLDDT of the PDB data where the pLDDT score of PDB is similar to DPLM, suggesting that DPLM learns to generate protein sequences with overall similar structural characteristics as PDB, as shown in Tab. 3.
Table 3: pLDDT score of UniRef50, PDB and DPLM unconditional sampling.
Length UniRef50 PDB DPLM
100 66.54 84.62 70.66
200 68.61 79.32 83.55
300 78.30 84.51 82.39
400 79.80 80.49 86.75
500 75.17 77.79 82.56
avg. pLDDT 73.68 81.34 81.18
Regarding mode collapse.
We also want to investigate whether DPLM collapses into the modes with high pLDDT sequences. To verify this, we evaluate the pseudo-perplexity of DPLM against subsets of UR50 sequences of high pLDDT and low pLDDT. The results are shown in Tab. 4. We can find that the ppl of less-structural proteins (pLDDT $<$ 50) is similar to the structural proteins (pLDDT $>$ 70), suggesting that DPLM equally learns protein sequences with diverse structural patterns.
Table 4: pseudo-ppl of less structural and more structural sequences.
less structural sequences more structural sequences
pseudo-ppl 2.36 2.55
In conclusion, we would like to provide a possible explanation for this phenomenon. We suggest that from a perspective of probabilistic graphical model (PGM), more structured data is generally more easy to learn due to stronger correlation between its elements. As such, we hypothesize the learning dynamics of protein LMs from through evolutionary sequences is first to digest those more structural proteins as co-evolutionary effects between amino acids play a prevailing and vital role in folding patterns, and then start to learn those less structural folding patterns, which could be long-tailed.
This could somehow relate to the so-called emergence phenomenon in the realm of LLMs, where scaling up LLMs leads to ”grokking” those long-tailed abilities. We would leave a study of the learning dynamics of protein LM as an exciting future investigation and hopefully can bring some interesting insights to the community.
D.3 Sequence-conditional generation: motif-scaffolding
@@
D.4 Structure-conditional generation: inverse folding
@@
$\mathcal{J}_{t}=\mathbb{E}_{q(\mathbf{x}^{(0)})}\bigg{[}\lambda^{(t)}\sum_{1\leq i\leq L}b_{i}(t)\cdot\log p_{\theta}(\mathbf{x}^{(0)}_{i}|\mathbf{x}^{(t)}_{\textrm{draft}})\bigg{]},$ (7)
@@
D.5 Classifier-free guidance
@@
Appendix E Related Work
E.1 Language Models
The dominant paradigm of language models is autoregressive language models, which breaks down the mutual distribution over the tokens of a sequence into conditional probabilities via the chain rule ${p(\mathbf{x}^{[1:N]})=\prod_{i=1}^{N}p(x^{[i]}|\mathbf{x}^{[1:i-1]})}$ and generates tokens by ancestral sampling from left to right (Bengio et al., 2000; Sutskever et al., 2014; Vaswani et al., 2017). Recently, researchers propose the non-autoregressive language models as an alternative (Gu et al., 2018). These models do not need to obey the left to right generation order (Qian et al., 2022; Huang et al., 2023) and demonstrate competitive or superior performance compared to their autoregressive counterpart across a wide range of domains including languages (Qian et al., 2021b; Huang et al., 2023; Qian et al., 2022; Huang et al., 2023; Zheng et al., 2023a), speeches (Kim et al., 2021), proteins (Zheng et al., 2023b), and molecules (Hoogeboom et al., 2022). Among the numerous non-autoregressive language models, diffusion language models (Li et al., 2022; Gong et al., 2022; Zheng et al., 2023a) have emerged as a solid and promising framework. Pretraining language models on a massive scale of unlabeled data markedly improves their downstream task performance (Mikolov et al., 2013; Peters et al., 2018; Radford et al., 2018; Devlin et al., 2019). As data volume and model sizes scale up, the training loss of language models predictably declines (Kaplan et al., 2020; Hoffmann et al., 2022; Muennighoff et al., 2023), and enhancing downstream task performance even without specific tuning (Radford et al., 2019). GPT3 (Brown et al., 2020) is a significant point in the journey, taking model sizes to 175B parameters, proposing in-context learning to bolster language models’ competence in solving certain tasks with only a handful of demonstrations. Furthermore, Wei et al. (2021); Sanh et al. (2022); Ouyang et al. (2022) introduce instruction tuning, finetuning pretrained language models on series of tasks described via instructions, which elicits the instruction following ability of models and significantly enhances their zero-shot performance on unseen tasks. More impressively, sufficiently large language models exhibit the emergent abilities such as multi-step reasoning (Kojima et al., 2022; Wei et al., 2022a, b), which small models do not possess (Fu et al., 2023). Empowered by large language models, helpful applications such as conversational AI systems [^1] https://chat.openai.com/ and autonomous agents [^2] https://github.com/Significant-Gravitas/Auto-GPT have garnered much interest. Although the most capable models at the moment are restricted in access, open-sourced efforts (Zeng et al., 2022; Touvron et al., 2023a, b; Taori et al., 2023; Chiang et al., 2023; Sun & Qiu, 2023) have largely enhanced the public accessibility of powerful large language models.
E.2 Protein Language Models
@@
E.3 Diffusion Language Models
@@
E.4 Protein Structure Diffusion Models
@@
E.5 Protein Inverse Folding
@@
Appendix F Visualization of Unconditional Samples